# Understanding Multimodal LLMs Under Distribution Shifts: An Information-Theoretic Approach

### v1

@@ 7 Discussion
This work urged the development of a formal framework to understand MLLMs under distribution shifts which is unexplored yet crucial for reliable artificial intelligence in the wild. As a first step for this, we devised effective mutual information (EMI) as a metric for MLLM evaluation and showed its theoretical connection to an existing standard metric, win rate. Then, we provide a theoretical upper bound for an MLLM’s EMI difference between ID and OOD that consists of JS divergence terms for marginal and conditional distributions of input/output variables. Through experiments on a benchmark spanning 61 distribution shifts, we show the correlation between win rate and EMI, and further show the correlation between EMI difference and its upper bound thereby empirically verifying our theoretical upper bound.
@@
As shown in Table 2, EMI strongly correlates with win rate. Compared to the win rate, the MI estimator can be computed more efficiently without relying on computationally expensive judge LLM (Achiam et al., 2023) (see Appendix C.4 for details). Therefore, EMI can be leveraged as a cost-efficient and theoretically grounded evaluation metric that measures effective relevance between multimodal queries and open-ended responses. Besides, the upper bound of EMID can be adopted as a regularizer during post-training or test-time adaptation of MLLM to improve its robustness to distribution shifts (Li et al., 2023a).
@@
Although input-output relevance measured by EMI is one of the most important properties for instruction-following assistant models, other crucial quality attributes are not captured by the form of relevance term. Extending the theory to support evaluation across multiple facets of MLLM will be promising future work direction. Besides, we have simulated some intuitive types of distribution shifts with the simplified assumption for data structure, while leaving some complex shifts driven by spurious correlation (Simon, 1954) that may be covered by Theorem 4.6. Validation of theorems on such non-trivial distribution shifts could be an important extension.
@@
This work lays the first theoretical foundation for quantifying the reliability of MLLMs. The theoretical statements we made shine the light on analyzing the MLLM performance gap under distribution shifts which commonly emerged in real-world applications. Our framework can help aware of the potential risk, i.e., performance variation, of MLLMs, and guide the practitioners to devise a method towards robust adaptation for chat assistants, thereby ensuring the trustworthiness of artificial intelligence solutions in crucial social applications such as finance and healthcare.
@@
Appendix A Additional Description for Effective Mutual Information
@@
In Section 4.2, we provide some justifications for using EMI as an evaluation metric of MLLMs by revealing analogies to excess risk and effective robustness and presenting its theoretical connection to win rate. While LLM-as-a-Judge enables flexible evaluation for open-ended generation tasks with multiple user-defined criteria, EMI confines the facet of evaluation to query-response relevance. However, compromise in the flexibility of evaluation endows us to build solid theoretical statements that are necessary for understanding MLLMs and improving them in a principled way.
@@ Appendix B Implementation Details
Following (Cheng et al., 2020, 2021), we parameterize the $q_{\psi}$ as a multi-variate Gaussian distribution and estimate the mean and variance parameters of Gaussian with separated two-layer MLPs with 250 hidden dimension size. During mini-batch training, those MLPs consume the concatenated input and response embeddings $\{[z_{\mathbf{x}_{i}},z_{y_{i}}]\}_{i=1}^{N}$ to produce a scalar estimate of MI, and they are simultaneously optimized by AdamW optimizer with learning rate 0.001 and batch size 1,024 for 5,000 iterations. However, if we have to train an estimator for every ID-OOD data pair, it may not be practical when the number of data pairs to be evaluated is large. Therefore, we constructed a dataset that integrates all ID-OOD data subsets for MI training (integration of all variants of LLaVA-Bench datasets reach roughly 5,000 samples for natural shift, and 9,000 samples for synthetic shift), trains it only once, and then infers all ID-OOD scenarios (27 for natural shift, 34 for synthetic shift) using this common MI estimator. This not only significantly reduces the time required to evaluate the model’s robustness against multiple distribution shift scenarios, but also stabilizes the training process by increasing the size of the data set used in the training process.
@@
MLLM judge and win rate.
For the open-ended generation tasks, (M)LLM-as-a-Judge has been adopted as a current de facto standard. Following (Liu et al., 2023, 2024a), we use GPT-4333gpt-4-turbo with 2024-08-01-preview API version was adopted with text-only inference mode (with plain-text form visual cue such as ground truth caption for image) as a judge model and also use the output of the same model as a reference answer for each query. We leverage the prompts provided by the source code of LLaVA [^4] https://github.com/haotian-liu/LLaVA, and compute the win rate of a model of interest by comparing its output with that of GPT-4.
@@ C.1 Additional Result from Pilot Study
In Section 3, we conduct an experiment to validate our hypotheses on the relation between MLLM performance degradation and the severity of natural distribution shift. In Figure 6, we provide additional results from another type of distribution shift that occurred by image and text perturbations. For image perturbation, we consider defocus blur and frost with three different magnitudes, and for text perturbation, we consider keyboard typo error and word synonym replacement with two different magnitudes. We observe the consistent trend in the relation between MLLM performance degradation and the severity of distribution shifts for the case of visual-only, text-only, and joint shift, likewise the case of natural shifts in Figure 1. That is, the increased magnitude of distribution shifts induces more severe MLLM performance degradation, and the degree of performance degradation can attribute to shifts in two modalities.
@@ C.2 Different Design Choices of MI and JSD Estimation
Table 4: Ablation study for MI estimator and embedding space. We evaluate four MI estimators with three different embedding space choices in terms of Spearman correlation coefficient $\rho$ between win rate and EMI. We can see that EMI and win rate are robustly correlated to variations in the embedding space and the MI estimator, but CLUB shows the most stable correlation.
@@
Table 5: Ablation study for MI estimator, JSD estimator, and embedding space. We evaluate four MI estimator and two JSD estimator candidates, with three different embedding space choices in terms of Pearson correlation coefficient between EMID and its upper bound. In all the considered variations, EMID and the upper bound of EMID (i.e., the simplified version in Theorem 4.5) show strong correlations, implying that our theorem empirical robustly holds in practice.
@@
In Table 4, we conduct Spearman correlation analysis over the 12 (4 $\times$ 3) cases of MI estimator and embedding space ablation. We can clearly see that EMI is consistently correlated with the win rate which demonstrates the robust effectiveness of our theorem in practice. Meanwhile, among the candidate MI estimators and embedding space, CLUB and two E5-V joint embeddings show outstanding results. However, E5-V embedding extraction require a forward pass of MLLM in contrast to the case of leveraging relatively small individual models (ViT base and BERT-base). To strike the balance between effectiveness and efficiency, we adopt CLIP ViT-B/32 and XLM-RoBERTa-Base embedding spaces by default.
@@ C.3 Hyperparameter Sensitivity
Figure 7: Sensitivity analysis for batch size, learning rate, and hidden dimension during MI estimator training. For the considered hyperparameter searching grid, MI estimates derived by the CLUB estimator robustly achieve high Spearman correlation with win rate (the largest deviation is less than 0.02 absolute value).
@@ C.4 Runtime Analysis
In addition to the advantage of allowing rigorous theoretical statements, EMI also has practical advantages over the win rate derived by the LLM judge. Specifically, while both win rate and EMI are model-dependent, the former relies on models with tens to hundreds of billions of parameters, while the latter enables meaningful evaluation even with relatively small embedding models with millions of parameters. To quantitatively argue this, we compare the time per instance and the total inference time for the entire LLaVA-Bench COCO dataset in Table 7. As shown in the table, EMI can shorten the time by 138 times compared to the win rate. Even including the time required for MI training, EMI-based evaluation is still 3 times faster than MLLM judgment-based evaluation. Since MI training is performed only once and then transferred to all ID-OOD scenarios, the efficiency of EMI-based evaluation over MLLM judgment becomes more evident as the number of datasets to be evaluated increases. In addition, in the LLM judge paradigm, although open-source LLM judges (Kim et al., 2024) have been actively studied recently, proprietary LLMs are still dominant in practice, so one must pay per instance query, whereas EMI can make meaningful inferences with publicly available open-source pre-trained models without paying per query.
Appendix D Proof and Additional Theorem
@@ D.1 Proof for the relationship between EMI and preference model
$|\text{EMI}(P_{\mathbf{X}Y};\theta)-\text{PM}(P_{\mathbf{X}Y};\theta)|\leq\delta+4.4\delta^{\frac{1}{8}}.$
@@
$\text{EMI}(P_{\mathbf{X}Y};\theta)$ $=I(P_{\mathbf{X}}\otimes P_{\theta})-I(P_{\mathbf{X}Y})$
@@
$\text{PM}(P_{\mathbf{X}Y};\theta)=\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[\mathbb{E}_{\hat{y}\sim P_{\theta}(\cdot|\mathbf{x})}\log P_{Y|\mathbf{X}=\mathbf{x}}(\hat{y})-\mathbb{E}_{y\sim P_{Y|\mathbf{X}=\mathbf{x}}}\log P_{Y|\mathbf{X}=\mathbf{x}}(y)]$
@@
$|\text{EMI}(P_{\mathbf{X}Y};\theta)-\text{PM}(P_{\mathbf{X}Y};\theta)|$ $=|H(P_{Y_{\theta}})-H(P_{Y})+\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[\mathbb{E}_{\hat{y}\sim P_{\theta}(\cdot|\mathbf{x})}\log\frac{P_{\theta}(\hat{y}|\mathbf{x})}{P_{Y|\mathbf{X}=\mathbf{x}}(\hat{y})}]|$
@@ D.2 Proof for EMID upper bound
where $P_{Y_{\theta}}=\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}P_{{\theta}}(\cdot|\mathbf{x})$ and $Q_{Y_{\theta}}=\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}P_{{\theta}}(\cdot|\mathbf{x})$, then $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta)$ is upper bounded by
@@
$\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta)$
$=\text{EMI}(P_{\mathbf{X}Y};\theta)-\text{EMI}(Q_{\mathbf{X}Y};\theta)$
$=(H(P_{Y_{\theta}})-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-H(P_{Y})+H(P_{Y|\mathbf{X}}))-((H(Q_{Y_{\theta}})-\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-H(Q_{Y})+H(Q_{Y|\mathbf{X}}))$
$\leq(H(P_{Y|\mathbf{X}})-H(Q_{Y|\mathbf{X}}))+\big{(}\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]\big{)}+|H(P_{Y_{\theta}})-H(P_{Y})+H(Q_{Y})-H(Q_{Y_{\theta}})|$
$\leq(H(P_{Y|\mathbf{X}})-H(Q_{Y|\mathbf{X}}))+\big{(}\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]\big{)}+|H(P_{Y_{\theta}})-H(P_{Y})|+|H(Q_{Y})-H(Q_{Y_{\theta}})|$
@@
$\small\begin{split}&=2D_{\rm JS}(P_{X_{v}X_{t}}||Q_{X_{v}X_{t}})\\ &=D_{\rm KL}(P_{X_{v}X_{t}}||M_{X_{v}X_{t}})+D_{\rm KL}(Q_{X_{v}X_{t}}||M_{X_{v}X_{t}})\\ &=\frac{1}{2}\big{(}D_{\rm KL}(P_{X_{v}}||M_{X_{v}})+\mathbb{E}_{P_{X_{v}}}D_{\rm KL}(P_{X_{t}|X_{v}=x_{v}}||M_{X_{t}|X_{v}=x_{v}})\big{)}+\frac{1}{2}\big{(}D_{\rm KL}(Q_{X_{v}}||M_{X_{v}})+\mathbb{E}_{Q_{X_{v}}}D_{\rm KL}(Q_{X_{t}|X_{v}=x_{v}}||M_{X_{t}|X_{v}=x_{v}})\big{)}\\ &+\frac{1}{2}\big{(}D_{\rm KL}(P_{X_{t}}||M_{X_{t}})+\mathbb{E}_{P_{X_{t}}}D_{\rm KL}(P_{X_{v}|X_{t}=x_{t}}||M_{X_{v}|X_{t}=x_{t}})\big{)}+\frac{1}{2}\big{(}D_{\rm KL}(Q_{X_{t}}||P_{X_{t}})+\mathbb{E}_{Q_{X_{t}}}D_{\rm KL}(Q_{X_{v}|X_{t}=x_{t}}||M_{X_{v}|X_{t}=x_{t}})\big{)}\\ &=D_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D_{\rm JS}(P_{X_{t}}||Q_{X_{t}})\\ &+\frac{1}{2}\big{(}\mathbb{E}_{P_{X_{v}}}D_{\rm KL}(P_{X_{t}|X_{v}=x_{v}}||M_{X_{t}|X_{v}=x_{v}})+\mathbb{E}_{Q_{X_{v}}}D_{\rm KL}(Q_{X_{t}|X_{v}=x_{v}}||M_{X_{t}|X_{v}=x_{v}})\big{)}\\ &+\frac{1}{2}\big{(}\mathbb{E}_{P_{X_{t}}}D_{\rm KL}(P_{X_{v}|X_{t}=x_{t}}||M_{X_{v}|X_{t}=x_{t}})+\mathbb{E}_{Q_{X_{t}}}D_{\rm KL}(Q_{X_{v}|X_{t}=x_{t}}||M_{X_{v}|X_{t}=x_{t}})\big{)}\\ &\leq D_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D_{\rm JS}(P_{X_{t}}||Q_{X_{t}})+\bar{D}_{\rm JS}(P_{X_{t}|X_{v}}||Q_{X_{t}|X_{v}})+\bar{D}_{\rm JS}(P_{X_{v}|X_{t}}||Q_{X_{v}|X_{t}})\end{split}$

### v2

@@ 7 Discussion
This work urged the development of a formal framework to understand MLLMs under distribution shifts which is unexplored yet crucial for reliable artificial intelligence in the wild. As a first step for this, we devised effective mutual information (EMI) as a metric for MLLM evaluation and showed its theoretical connection to an existing standard metric, relative preference score. Then, we provide a theoretical upper bound for an MLLM’s EMI difference between ID and OOD that consists of JS divergences for marginal and conditional distributions of input/output variables. Through experiments on benchmarks spanning 61 distribution shifts, we show the correlation between relative preference and EMI, and further show correlations between EMI difference and its upper bound, thereby empirically verifying our theoretical claims across various models.
@@
As shown in Table 2, EMI strongly correlates with RP. Compared to RP, the MI estimator can be computed more efficiently without relying on the computationally expensive judge LLM (Achiam et al., 2023) (see Appendix C.4). Thus, EMI can be used as a theoretically grounded, cost-effective evaluation metric that measures the effective relevance between multimodal queries and open-ended responses. Besides, the upper bound of EMID can be adopted as a regularizer during model training, as we showed in Table 6, or test-time adaptation of MLLM to improve its robustness to distribution shifts (Li et al., 2023a), as well as a robustness measure of MLLM.
@@
Although input-output relevance measured by EMI is one of the most important properties for instruction-following models, other crucial quality attributes are not captured by the form of the relevance term. Extending the theory to support evaluation across multiple facets of MLLM will be worth exploring. In addition, we only simulated some intuitive types of distribution shifts with a simplified assumption for the data structure, leaving the analysis on some complex shifts driven by spurious correlation (Simon, 1954) that may be covered by Thm. 4.6. Meanwhile, despite its consistent correlation to EMID, our upper bound is not quite tight in theory, and we did not discuss a lower bound of EMID. Pursuing a tighter bound or exploring the lower limit of EMID can also be worthwhile.
Acknowledgement
We gratefully appreciate the ICML anonymous reviewers for their valuable feedback, and also appreciate constructive comments from Max Khanov, Min-Hsuan Yeh, Dongkwan Kim, and Kyungeun Lee. Changdae Oh, Shawn Im, Xuefeng Du, and Yixuan Li are supported by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation (NSF) Award No. IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, and Alfred P. Sloan Fellowship. Also, Zhen Fang was funded by the Australian Government through the Australian Research Council (ARC) under grant number DE250100363.
@@ Acknowledgement
This work lays the first theoretical foundation for quantifying the reliability of MLLMs. The theoretical statements we made shine a light on analyzing the MLLM performance gap under distribution shifts, which commonly emerged in real-world applications. Our framework can help aware of the potential risk, i.e., performance variation, of MLLMs, and guide practitioners to devise a method towards robust adaptation for chat assistants, thereby ensuring the trustworthiness of artificial intelligence solutions in crucial social applications such as finance and healthcare.
@@
Appendix A Detailed Description for Effective Mutual Information
@@
In Section 4.2, we provide some justifications for using EMI as an evaluation metric of MLLMs by revealing analogies to excess risk and effective robustness and presenting its theoretical connection to relative preference score. While LLM-as-a-Judge enables flexible evaluation for open-ended generation tasks with multiple user-defined criteria, EMI confines the facet of evaluation to query-response relevance. However, compromise in the flexibility of evaluation enables us to build solid theoretical statements that are necessary for understanding MLLMs to shifts and improving them in a principled way.
@@ Appendix B Implementation Details
Following (Cheng et al., 2020, 2021), we parameterize the $q_{\psi}$ as a multi-variate Gaussian distribution and estimate the mean and variance parameters of Gaussian with separated two-layer MLPs with 250 hidden dimension size. During mini-batch training, those MLPs consume the concatenated input and response embeddings $\{[z_{\mathbf{x}_{i}},z_{y_{i}}]\}_{i=1}^{N}$ to produce a scalar estimate of MI, and they are simultaneously optimized by AdamW optimizer with learning rate 0.001 and batch size 1,024 for 5,000 iterations. However, if we have to train an estimator for every ID-OOD data pair, it may not be practical when the number of data pairs to be evaluated is large. Therefore, we constructed a dataset that integrates all ID-OOD data subsets for MI training (integration of all variants of LLaVA-Bench datasets reach roughly 5,000 samples for natural shift, and 9,000 samples for synthetic shift), trains it only once, and then infers all ID-OOD scenarios (27 for natural shift, 34 for synthetic shift) using these common MI estimators. This not only significantly reduces the time required to evaluate the model’s robustness against multiple distribution shift scenarios, but also stabilizes the training process by increasing the size of the data set used in the training process.
@@
Scale-adjusted EMID upper bound construction.
By leveraging the estimator described above, one can compute all JSD terms in the proposed EMID upper bound (Eq. 10): $D_{\rm JS}(P_{X_{v}}||Q_{X_{v}})$, $D_{\rm JS}(P_{X_{t}}||Q_{X_{t}})$, $D_{\rm JS}(P_{Y_{\theta}}||P_{Y})$, and $D_{\rm JS}(Q_{Y_{\theta}}||Q_{Y})$). Because the exact computation of the entropy scaler term $\hat{H}$ in Eq. (10) is intractable, we relax it with an estimate on batch samples, and replace the inaccessible true conditional distribution $Q_{Y|\mathbf{X}}$ with $P_{\theta}$. For a case of sentence output $Y=\{y_{1},...,y_{L}\}$ with $L$ tokens, the length-normalized batch entropy estimate over $N$ samples indexed through $i\in\mathcal{I}=\{1,...,N\}$ can be formulated as below,
$\tilde{H}=\max_{i\in\mathcal{I}}-\frac{1}{L}\sum_{l=1}^{L}\log P_{\theta}(y_{i,l}|y_{i,<l},x_{i}).$ (14)
In pilot experiments, we observed that the values of $\tilde{H}$ are centered on 2.0 in the datasets that we considered in this work. Therefore, $\hat{H}=\max_{\mathbf{x}\in\mathcal{X}}[H(Q_{Y|\mathbf{X}=\mathbf{x}})+H(P_{\theta}(\cdot|\mathbf{x}))]$ can approximately have a value 4.0. As the main implications of EMID upper bound are providing the characterization of MLLM performance gaps and serving as a practical measure of robustness, we compute the scale-adjusted upper bound estimates for EMID upper bound (UB) as below,
EMID UB $=\widehat{H}\big{(}{D^{\frac{1}{2}}_{\rm JS}(P_{X_{v}}\|Q_{X_{v}})}+{D^{\frac{1}{2}}_{\rm JS}(P_{X_{t}}\|Q_{X_{t}})}\big{)}+4(D^{\frac{1}{4}}_{\rm JS}(P_{Y_{\theta}}||P_{Y})+D^{\frac{1}{4}}_{\rm JS}(Q_{Y_{\theta}}||Q_{Y}))$
$\approx 4\big{(}{D^{\frac{1}{2}}_{\rm JS}(P_{X_{v}}\|Q_{X_{v}})}+{D^{\frac{1}{2}}_{\rm JS}(P_{X_{t}}\|Q_{X_{t}})}+D^{\frac{1}{4}}_{\rm JS}(P_{Y_{\theta}}||P_{Y})+D^{\frac{1}{4}}_{\rm JS}(Q_{Y_{\theta}}||Q_{Y})\big{)}$
[…]
See Section D for the detailed derivation of EMID UB. Note that applying the linear transformation to the target variables of Pearson correlation analysis does not affect the correlation coefficient value, so that the correlation coefficients between the EMID and its scale-adjusted UB has the same values as between the EMID and its original UB, even though it is not the exact estimate of EMID UB.
@@
MLLM judge and relative preference (RP) score.
For evaluation of open-ended generation tasks, (M)LLM-as-a-Judge has been adopted as a current de facto standard. Following (Liu et al., 2023, 2024a), we use GPT-4333gpt-4-turbo with 2024-08-01-preview API version was mainly adopted with text-only inference mode (with plain-text form visual cue such as ground truth caption for image) as a judge model and also use the output of the same model as a reference answer for each query. We leverage the prompts provided by the source code of LLaVA [^4] https://github.com/haotian-liu/LLaVA, and compute the RP score of a model of interest by comparing its output with that of the reference answer.
@@ C.1 Additional Result from Pilot Study
In Section 3, we conduct an experiment to validate our hypotheses on the relation between MLLM performance degradation and the severity of natural distribution shift. In Figure 6, we provide additional results from another type of distribution shift that occurred by image and text perturbations. For image perturbation, we consider defocus blur and frost with three different magnitudes, and for text perturbation, we consider keyboard typo error and word synonym replacement with two different magnitudes (We adopt the source code of https://github.com/Jielin-Qiu/MM_Robustness to generate perturbed datasets). We observe the consistent trend in the relation between MLLM performance degradation and the severity of distribution shifts for the case of visual-only, text-only, and joint shift, likewise the case of natural shifts in Figure 1. That is, the increased magnitude of distribution shifts induces more severe MLLM performance degradation, and the degree of performance degradation can attribute to shifts in two modalities.
@@ C.2 Different Design Choices of MI and JSD Estimation
Table 7: Ablation study for MI estimator and embedding space. We evaluate four MI estimators with three different embedding space choices in terms of Spearman correlation coefficient $\rho$ between RP score and EMI. We can see that EMI and RP score are robustly correlated to variations in the embedding space and the MI estimator, but CLUB shows the most stable correlation.
@@
Table 8: Ablation study for MI estimator, JSD estimator, and embedding space. We evaluate four MI estimator candidates and two JSD estimator candidates, with three different embedding space choices in terms of Pearson correlation coefficient between EMID and its upper bound. In all the considered variations, EMID and the upper bound of EMID (i.e., the simplified version in Theorem 4.5) show strong correlations, implying that our theorem robustly holds in practice.
@@
In Table 7, we conduct Spearman correlation analysis over the 12 (4 $\times$ 3) cases of MI estimator and embedding space ablation. We can clearly see that EMI is consistently correlated with the RP score which demonstrates the robust effectiveness of our theorem in practice. Meanwhile, among the candidate MI estimators and embedding space, CLUB and two E5-V joint embeddings show outstanding results. However, E5-V embedding extraction require a forward pass of MLLM in contrast to the case of leveraging relatively small individual models (ViT base and BERT-base). To strike the balance between effectiveness and efficiency, we adopt CLIP ViT-B/32 and XLM-RoBERTa-Base embedding spaces by default.
@@ C.3 Hyperparameter Sensitivity
Figure 7: Sensitivity analysis for batch size, learning rate, and hidden dimension during MI estimator training. For the considered hyperparameter searching grid, MI estimates derived by the CLUB estimator robustly achieve high Spearman correlation with RP score (the largest deviation is less than 0.02 absolute value).
@@ C.4 Runtime Analysis
In addition to the advantage of allowing rigorous theoretical statements, EMI also has practical advantages over the RP score derived by the LLM judge. Specifically, while both RP score and EMI are model-dependent, the former relies on models with tens to hundreds of billions of parameters, while the latter enables meaningful evaluation even with relatively small embedding models with millions of parameters. To quantitatively argue this, we compare the inference time per instance and for the entire LLaVA-Bench COCO dataset in Table 10. As shown in the table, EMI can shorten the time by 138 times compared to the RP score. Even including the time required for MI training, EMI-based evaluation is still 3 times faster than MLLM judgment-based evaluation. Since MI training is performed only once and then transferred to all ID-OOD scenarios, the efficiency of EMI-based evaluation over MLLM judgment becomes more evident as the number of datasets to be evaluated increases. In addition, in the LLM judge paradigm, although open-source LLM judges (Kim et al., 2024) have been actively studied recently, proprietary LLMs are still dominant in practice, so one must pay per instance query, whereas EMI can make meaningful inferences with publicly available open-source feature extractors without paying per query.
C.5 Notes on Practical Usage
Although the empirical estimates of EMI and EMID upper bound show consistent correlations with the relative preference score and EMID, the absolute value of MI estimates varies depending on the estimator types and their configurations. Therefore, if one has used the empirical EMI estimates (computed by a two-layer MLP-based CLUB estimator on CLIP-VIT and RoBERTa embedding spaces) to assess the qualities of responses from MLLMs, the same estimator and the same embedding models should be used across targeted models and datasets for fair comparison.
Appendix D Extended Theoretical Analysis with Full Proof
@@ D.1 Proof for the relationship between EMI and preference model
$|\text{EMI}(P_{\mathbf{X}Y};P_{\theta})-\text{PM}(P_{\mathbf{X}Y};P_{\theta})|\leq\delta+4.4\delta^{\frac{1}{8}}.$
@@
$\text{EMI}(P_{\mathbf{X}Y};P_{\theta})$ $=I(P_{\mathbf{X}}\otimes P_{\theta})-I(P_{\mathbf{X}Y})$
@@
$\text{PM}(P_{\mathbf{X}Y};P_{\theta})=\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[\mathbb{E}_{\hat{y}\sim P_{\theta}(\cdot|\mathbf{x})}\log P_{Y|\mathbf{X}=\mathbf{x}}(\hat{y})-\mathbb{E}_{y\sim P_{Y|\mathbf{X}=\mathbf{x}}}\log P_{Y|\mathbf{X}=\mathbf{x}}(y)]$
@@
$|\text{EMI}(P_{\mathbf{X}Y};P_{\theta})-\text{PM}(P_{\mathbf{X}Y};P_{\theta})|$ $=|H(P_{Y_{\theta}})-H(P_{Y})+\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[\mathbb{E}_{\hat{y}\sim P_{\theta}(\cdot|\mathbf{x})}\log\frac{P_{\theta}(\hat{y}|\mathbf{x})}{P_{Y|\mathbf{X}=\mathbf{x}}(\hat{y})}]|$
@@ D.2 Proof for EMID upper bound
where $P_{Y_{\theta}}=\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}P_{{\theta}}(\cdot|\mathbf{x})$ and $Q_{Y_{\theta}}=\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}P_{{\theta}}(\cdot|\mathbf{x})$, then $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};P_{\theta})$ is upper bounded by
@@
$\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};P_{\theta})$
$=\text{EMI}(P_{\mathbf{X}Y};P_{\theta})-\text{EMI}(Q_{\mathbf{X}Y};P_{\theta})$
$=(H(P_{Y_{\theta}})-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-H(P_{Y})+H(P_{Y|\mathbf{X}}))$
$-((H(Q_{Y_{\theta}})-\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-H(Q_{Y})+H(Q_{Y|\mathbf{X}}))$
$\leq(H(P_{Y|\mathbf{X}})-H(Q_{Y|\mathbf{X}}))+\big{(}\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]\big{)}$
$+|H(P_{Y_{\theta}})-H(P_{Y})+H(Q_{Y})-H(Q_{Y_{\theta}})|$
$\leq(H(P_{Y|\mathbf{X}})-H(Q_{Y|\mathbf{X}}))+\big{(}\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]\big{)}$
$+|H(P_{Y_{\theta}})-H(P_{Y})|+|H(Q_{Y})-H(Q_{Y_{\theta}})|$
@@
$\small\begin{split}&2D_{\rm JS}(P_{X_{v}X_{t}}||Q_{X_{v}X_{t}})\\ &=D_{\rm KL}(P_{X_{v}X_{t}}||M_{X_{v}X_{t}})+D_{\rm KL}(Q_{X_{v}X_{t}}||M_{X_{v}X_{t}})\\ &=\frac{1}{2}\big{(}D_{\rm KL}(P_{X_{v}}||M_{X_{v}})+\mathbb{E}_{P_{X_{v}}}D_{\rm KL}(P_{X_{t}|X_{v}=x_{v}}||M_{X_{t}|X_{v}=x_{v}})\big{)}\\ &+\frac{1}{2}\big{(}D_{\rm KL}(Q_{X_{v}}||M_{X_{v}})+\mathbb{E}_{Q_{X_{v}}}D_{\rm KL}(Q_{X_{t}|X_{v}=x_{v}}||M_{X_{t}|X_{v}=x_{v}})\big{)}\\ &+\frac{1}{2}\big{(}D_{\rm KL}(P_{X_{t}}||M_{X_{t}})+\mathbb{E}_{P_{X_{t}}}D_{\rm KL}(P_{X_{v}|X_{t}=x_{t}}||M_{X_{v}|X_{t}=x_{t}})\big{)}\\ &+\frac{1}{2}\big{(}D_{\rm KL}(Q_{X_{t}}||P_{X_{t}})+\mathbb{E}_{Q_{X_{t}}}D_{\rm KL}(Q_{X_{v}|X_{t}=x_{t}}||M_{X_{v}|X_{t}=x_{t}})\big{)}\\ &=D_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D_{\rm JS}(P_{X_{t}}||Q_{X_{t}})\\ &+\frac{1}{2}\big{(}\mathbb{E}_{P_{X_{v}}}D_{\rm KL}(P_{X_{t}|X_{v}=x_{v}}||M_{X_{t}|X_{v}=x_{v}})+\mathbb{E}_{Q_{X_{v}}}D_{\rm KL}(Q_{X_{t}|X_{v}=x_{v}}||M_{X_{t}|X_{v}=x_{v}})\big{)}\\ &+\frac{1}{2}\big{(}\mathbb{E}_{P_{X_{t}}}D_{\rm KL}(P_{X_{v}|X_{t}=x_{t}}||M_{X_{v}|X_{t}=x_{t}})+\mathbb{E}_{Q_{X_{t}}}D_{\rm KL}(Q_{X_{v}|X_{t}=x_{t}}||M_{X_{v}|X_{t}=x_{t}})\big{)}\\ &\leq D_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D_{\rm JS}(P_{X_{t}}||Q_{X_{t}})+\bar{D}_{\rm JS}(P_{X_{t}|X_{v}}||Q_{X_{t}|X_{v}})+\bar{D}_{\rm JS}(P_{X_{v}|X_{t}}||Q_{X_{v}|X_{t}})\end{split}$